# Hamiltonian Descent Algorithms for Optimization: Accelerated Rates via Randomized Integration Time

### v1

@@ G.3.3 Proof of Theorem 5 (convergence of RHGD under weak convexity)
Applying Propositions 2 and 1 with the choice $h\leq\frac{1}{8\sqrt{L}}\leq\frac{1}{4^{\nicefrac{{1}}{{8}}}\sqrt{L}}$, we have
$\frac{41(k+8)^{2}}{9}\|\mathcal{G}_{k+1}^{h}\|^{2}$ $\leq\frac{328(k+8)^{2}}{9}(1+L^{2}h^{4})\cdot L^{2}h^{8}\|\nabla f(x_{k+1})\|^{2}$
$\leq\frac{164(k+8)^{2}L^{2}h^{8}}{3}\|\nabla f(x_{k+1})\|^{2}$
$\leq\frac{41(k+8)^{2}h^{4}}{3072}\|\nabla f(x_{k+1})\|^{2}$
@@
where the second inequality follows from $h\leq\frac{1}{4^{\nicefrac{{1}}{{8}}}\sqrt{L}}$ and the third inequality follows from $h\leq\frac{1}{8\sqrt{L}}$. Combining with (36), we have
@@
where the third inequality follows from $L$-smoothness of $f$ and the last inequality follows from $h\leq\frac{1}{8\sqrt{L}}$. ∎
@@ G.3.4 Proof of Corollary 2
If we choose $h=\frac{1}{8\sqrt{L}}$, we obtain from Theorem 5
$\mathbb{E}[f(x_{K})-f(x^{*})]$ $\leq\frac{896L\cdot\mathbb{E}\left[\|x_{0}-x^{*}\|^{2}\right]}{(K+8)^{2}}.$
@@
$\mathbb{E}[f(x_{K})-f(x^{*})]\leq\frac{896L\cdot\mathbb{E}\left[\|x_{0}-x^{*}\|^{2}\right]}{(K+8)^{2}}\leq\frac{896L\cdot\mathbb{E}\left[\|x_{0}-x^{*}\|^{2}\right]}{K^{2}}\leq\varepsilon,$
@@
$K\geq\sqrt{\frac{896L\cdot\mathbb{E}\left[\|x_{0}-x^{*}\|^{2}\right]}{\varepsilon}}.$
@@ H.1.1 Comparison with baseline algorithms
Figure LABEL:fig:quadcomp-add shows the convergence results for setting (1) with exact $\alpha$ under condition numbers $\kappa\in\{10^{3},10^{5}\}$. RHGD consistently outperforms GD while being comparable to AGD and CAGD.
Figure LABEL:fig:quadcompmismu-add presents the results for setting (2), where $\hat{\alpha}\in\{0.1,1\}$ is overestimated. We observe that while the performance of AGD and CAGD degrade significantly when $\alpha$ is poorly estimated, RHGD maintains robustness across different values of $\alpha$, demonstrating its practical advantage under parameter uncertainty.
Figure LABEL:fig:quadcompC-add provides the results for setting (3), i.e., when $\alpha=0$ (weakly convex), with smoothness constants $L\in\{5\times 10^{3},5\times 10^{4}\}$. RHGD outperform AGD and CAGD in late iterations while being noticeably faster than GD.
@@ H.1.2 Refreshment behavior
In the strongly convex setting, we choose $\kappa=10^{7}$ with $L=500$, and $h=\sqrt{1/L}$. In the weakly convex setting, we choose $L=500$ and $h=\sqrt{1/L}$. Then we plot the objective value at each iteration versus the accumulated Poisson time $T_{k}$, thereby aligning our discrete-time algorithms with their continuous-time interpretations. We also overlay the actual refreshment events as markers on the plot, which allows us to visually compare the algorithm’s progress with the stochastic timing of velocity refreshment. These visualizations shown in Figure LABEL:fig:poisson_plot confirm that our refreshment mechanisms closely match the intended Poisson-driven behavior and provide an intuitive connection between the discrete-time algorithm and their continuous-time limit flow.
@@ H.2 Logistic loss minimization
In the main paper, we only report the results of $\alpha\in\{10^{-4},0\}$. Here we provide the remaining results corresponding to $\{10^{-3},10^{-5}\}$. The setting $\alpha=10^{-3}$ corresponds to a commonly used regularization level in practice, providing a well-conditioned objective. In contrast, $\alpha=10^{-5}$ yields a more ill-conditioned problem, which serves as a stress test for the algorithms but is less typical in real-world applications. Figure LABEL:fig:logisc-add presents the convergence behavior of GD, AGD, CAGD, and RHGD under both settings. We observe that RHGD consistently outperforms GD and remains competitive with AGD and CAGD across both values of $\alpha$. Notably, RHGD using $\gamma_{k}=2\sqrt{\alpha}$ consistently achieves faster convergence than $\gamma_{k}=\sqrt{\alpha}$. These results further confirm the robustness and practical efficiency of RHGD under varying degrees of strong convexity.
@@ H.3 Details of baseline algorithms
$\theta_{k}^{\prime}$ $\tanh\left(\sqrt{\alpha\eta}\,\tau_{k}\right)$ 00
@@
5:         $\eta_{k+1}=1.1\cdot\eta_{k}$, and $x_{k+1}=\tilde{x}_{k+1}$triangleright Increase stepsize and accept trial step
@@
7:         $\eta_{k+1}=0.6\cdot\eta_{k}$, and $x_{k+1}=x_{k}$triangleright Decrease stepsize and reject trial step
@@
5:         $\eta_{k+1}=1.1\cdot\eta_{k}$, and $x_{k+1}=\tilde{x}_{k+1}$triangleright Increase stepsize and accept trial step
@@
7:         $\eta_{k+1}=0.6\cdot\eta_{k}$, and $x_{k+1}=x_{k}$triangleright Decrease stepsize and reject trial step
@@
8:         $\eta_{k+1}=1.1\cdot\eta_{k}$, and $x_{k+1}=\tilde{x}_{k+1}$triangleright Increase stepsize and accept trial step
@@
10:         $\eta_{k+1}=0.6\cdot\eta_{k}$, and $x_{k+1}=x_{k}$triangleright Decrease stepsize and reject trial step
@@
6:         $h_{k+1}=\sqrt{1.1}\cdot h_{k}$, and $x_{k+1}=\tilde{x}_{k+1}$triangleright Increase stepsize and accept trial step
@@
8:         $h_{k+1}=\sqrt{0.6}\cdot h_{k}$, and $x_{k+1}=x_{k}$triangleright Decrease stepsize and reject trial step

### v2

@@ G.3.3 Proof of Theorem 5 (convergence of RHGD under weak convexity)
Applying Propositions 2 and 1 with the choice $h\leq\frac{1}{7\sqrt{L}}\leq\frac{1}{4^{\nicefrac{{1}}{{8}}}\sqrt{L}}$, we have
$\frac{41(k+8)^{2}}{9}\|\mathcal{G}_{k+1}^{h}\|^{2}$ $\leq\frac{164(k+8)^{2}}{9}\cdot(1+L^{2}h^{4})\cdot\frac{L^{2}h^{8}}{1-2L^{4}h^{8}}\cdot\|\nabla f(x_{k+1})\|^{2}$
$\leq\frac{82(k+8)^{2}h^{4}}{3}\cdot\frac{L^{2}h^{4}}{1-2L^{4}h^{8}}\cdot\|\nabla f(x_{k+1})\|^{2}$
@@
where the second inequality follows from $h\leq\frac{1}{4^{\nicefrac{{1}}{{8}}}\sqrt{L}}$ and the third inequality follows from $h\leq\frac{1}{7\sqrt{L}}$. Combining with (36), we have
@@
where the third inequality follows from $L$-smoothness of $f$ and the last inequality follows from $h\leq\frac{1}{7\sqrt{L}}$. ∎
@@ G.3.4 Proof of Corollary 2
If we choose $h=\frac{1}{7\sqrt{L}}$, we obtain from Theorem 5
$\mathbb{E}[f(x_{K})-f(x^{*})]$ $\leq\frac{686L\cdot\mathbb{E}\left[\|x_{0}-x^{*}\|^{2}\right]}{(K+8)^{2}}.$
@@
$\mathbb{E}[f(x_{K})-f(x^{*})]\leq\frac{686L\cdot\mathbb{E}\left[\|x_{0}-x^{*}\|^{2}\right]}{(K+8)^{2}}\leq\frac{686L\cdot\mathbb{E}\left[\|x_{0}-x^{*}\|^{2}\right]}{K^{2}}\leq\varepsilon,$
@@
$K\geq\sqrt{\frac{686L\cdot\mathbb{E}\left[\|x_{0}-x^{*}\|^{2}\right]}{\varepsilon}}.$
@@ H.1.1 Comparison with baseline algorithms
Figure 5(a) shows the convergence results for setting (1) with exact $\alpha$ under condition numbers $\kappa\in\{10^{3},10^{5}\}$. RHGD consistently outperforms GD while being comparable to AGD and CAGD.
(a) Comparison between GD, AGD, CAGD, and RHGD (ours) on minimizing quadratic functions with $\kappa\in\{10^{3},\,10^{5}\}$ and optimal stepsizes via grid search. All algorithms use exact $\alpha$. Each plot shows results averaged over 5 runs.
[…]
(b) Comparison between GD, AGDCAGD and RHGD (ours) on minimizing the quadratic function with with $\kappa=10^{7}$ and optimal stepsizes via grid search. All algorithms use misspecified $\hat{\alpha}\in\{0.1,\,1\}$. Each plot shows results averaged over 5 runs.
Figure 5(c) provides the results for setting (3), i.e., when $\alpha=0$ (weakly convex), with smoothness constants $L\in\{5\times 10^{3},5\times 10^{4}\}$. RHGD outperform AGD and CAGD in late iterations while being noticeably faster than GD.
(c) Comparison between GD, AGDCAGD and RHGD with their optimal stepsizes via grid search for minimizing quadratic functions with $\alpha=0$ and $L\in\{5\times 10^{3},5\times 10^{4}\}$. Each plot shows results averaged over 5 runs.
(d) Comparison of actual refresh events in RHGD with theoretical Poisson jump times. Left: $\gamma_{k}=\sqrt{\alpha}$ vs. homogeneous Poisson process with rate $\sqrt{\alpha}$. Right: $\gamma_{k}=\frac{17}{2(k+9)h}$ vs. inhomogeneous Poisson process with rate $\gamma(t)=\frac{17}{2t+18h}$. The $x$-axis shows time $t=kh$ with $h=\sqrt{0.002}$. Blue dots: actual refreshes; gray dashed lines: Poisson jumps.
@@ H.1.2 Refreshment behavior
In the strongly convex setting, we choose $\kappa=10^{7}$ with $L=500$, and $h=\sqrt{1/L}$. In the weakly convex setting, we choose $L=500$ and $h=\sqrt{1/L}$. Then we plot the objective value at each iteration versus the accumulated Poisson time $T_{k}$, thereby aligning our discrete-time algorithms with their continuous-time interpretations. We also overlay the actual refreshment events as markers on the plot, which allows us to visually compare the algorithm’s progress with the stochastic timing of velocity refreshment. These visualizations shown in Figure 5(d) confirm that our refreshment mechanisms closely match the intended Poisson-driven behavior and provide an intuitive connection between the discrete-time algorithm and their continuous-time limit flow.
@@ H.2 Logistic loss minimization
In the main paper, we only report the results of $\alpha\in\{10^{-4},0\}$. Here we provide the remaining results corresponding to $\{10^{-3},10^{-5}\}$. The setting $\alpha=10^{-3}$ corresponds to a commonly used regularization level in practice, providing a well-conditioned objective. In contrast, $\alpha=10^{-5}$ yields a more ill-conditioned problem, which serves as a stress test for the algorithms but is less typical in real-world applications. Figure 7(a) presents the convergence behavior of GD, AGD, CAGD, and RHGD under both settings. We observe that RHGD consistently outperforms GD and remains competitive with AGD and CAGD across both values of $\alpha$. Notably, RHGD using $\gamma_{k}=2\sqrt{\alpha}$ consistently achieves faster convergence than $\gamma_{k}=\sqrt{\alpha}$. These results further confirm the robustness and practical efficiency of RHGD under varying degrees of strong convexity.
(a) Comparison of GD, AGD, CAGD and RHGD on logistic regression with $\alpha\in\{10^{-3},10^{-5}\}$. We run each algorithm using adaptive stepsizes. For RHGD, we evaluate $\gamma_{k}=\sqrt{\alpha}$ and $\gamma_{k}=2\sqrt{\alpha}$.
@@ H.3 Details of baseline algorithms
$\theta_{k}^{\prime}$ $\tanh\left(\sqrt{\alpha\eta}\,\tau_{k}\right)$ 0
@@
5:   $\eta_{k+1}=1.1\cdot\eta_{k}$, and $x_{k+1}=\tilde{x}_{k+1}$ $\triangleright$ Increase stepsize and accept trial step
@@
7:   $\eta_{k+1}=0.6\cdot\eta_{k}$, and $x_{k+1}=x_{k}$ $\triangleright$ Decrease stepsize and reject trial step
@@
5:   $\eta_{k+1}=1.1\cdot\eta_{k}$, and $x_{k+1}=\tilde{x}_{k+1}$ $\triangleright$ Increase stepsize and accept trial step
@@
7:   $\eta_{k+1}=0.6\cdot\eta_{k}$, and $x_{k+1}=x_{k}$ $\triangleright$ Decrease stepsize and reject trial step
@@
8:   $\eta_{k+1}=1.1\cdot\eta_{k}$, and $x_{k+1}=\tilde{x}_{k+1}$ $\triangleright$ Increase stepsize and accept trial step
@@
10:   $\eta_{k+1}=0.6\cdot\eta_{k}$, and $x_{k+1}=x_{k}$ $\triangleright$ Decrease stepsize and reject trial step
@@
6:   $h_{k+1}=\sqrt{1.1}\cdot h_{k}$, and $x_{k+1}=\tilde{x}_{k+1}$ $\triangleright$ Increase stepsize and accept trial step
@@
8:   $h_{k+1}=\sqrt{0.6}\cdot h_{k}$, and $x_{k+1}=x_{k}$ $\triangleright$ Decrease stepsize and reject trial step